# Inhomogeneous mixing: From microscopic dynamics to mesoscopic staircases

### v1

@@ Abstract
Inhomogeneous mixing and the consequent mesoscopic layered structure have been observed in many physical systems, including magnetically confined fusion plasmas. Especially, in plasmas, mixing can be enhanced through turbulence spreading by intermittent coherent structures (blobs/voids), or suppressed due to the formation of transport barriers (sheared zonal flows). Interestingly, blobs/voids and zonal flows are not independent, and they can co-exist in a state of inhomogeneous mixing, often called the $E\times B$ staircase. In this paper, we first introduce recent experimental progress on the physics of blobs/voids: how turbulence spreading by blobs/voids occurs, the consequences of enhanced turbulence spreading for the power decay length, and the interaction between blobs/voids and zonal flows. Then, we provide a brief review of experimental results on staircases, or more generally layered mesoscopic transport barriers. Staircases are often elusive and different complementary methods have been utilized to identify them, but our understanding is still incomplete. This paper serves as an initial step toward applying insights gained from inhomogeneous mixing due to blobs/voids to the understanding of a staircase.
@@ 1 Introduction
Inhomogeneous mixing is a transport or relaxation process which is not uniform. So, in Figure 1, case (a) depicts homogeneous mixing, while case (b) shows inhomogeneous mixing. Note there, the mixing zones don’t overlap. As suggested by Figure 1, inhomogeneous mixing of a profile will quite naturally produce layered or staircase structures on mesoscales or macroscales. Such layered structure is the focus of this collection, and staircases in magnetized plasma are the foci of this paper. Inhomogeneous mixing and staircases form in many physical systems, including saltwater, via thermohaline convection, stirred stratified fluids, geophysical fluids (i.e. potential vorticity (PV) staircases), Cahn-Hilliard Navier-Stokes fluids via phase separation and magnetized plasma, the subject of this volume. Inhomogeneous mixing has a bi-stable or multi-stable, and multi-scale character, which is intrinsic to the macroscopic layering, i.e. different regions mix unequally. This paper discusses inhomogeneous mixing and staircase structure in confined magnetized plasma.
Figure 1: Schematic diagram of (a) homogeneous mixing and (b) inhomogeneous mixing.
Turbulence in plasma (at low $\beta$) consists of an ensemble of spatially localized convective cells, pinned to sites determined by the confining magnetic geometry. $k\cdot B=0$ resonant or rational surfaces are one such pinning site. Typically, cells have radial coherent length $\lambda_{\mathrm{corr}}$ that is $\lambda_{\mathrm{corr}}/L_{p}\ll 1$. Here, $L_{p}$ is a plasma profile scale length. At the edge, $\lambda_{\mathrm{corr}}/L_{p}\leq 1$. There are three obvious means by which mixing by plasma turbulence develops inhomogeneity.
@@
    Along with particles, heat, etc., the turbulence mixes plasma potential vorticity – and thus vorticity – so $\langle{\tilde{v}}_{r}\nabla^{2}\tilde{\phi}\rangle\neq 0$. The Taylor identity then gives $\langle{\tilde{v}}_{r}\nabla^{2}\tilde{\phi}\rangle=-\partial\langle{\tilde{v}}_{r}{\tilde{v}}_{\theta}\rangle/\partial r$, so a non-zero $E\times B$ flow Reynolds force develops, and drives sheared zonal flows on the scale of the spectral envelope. These zonal flows sink energy from the underlying relaxation process and thus impose a variable shearing pattern on the turbulence. Consequently, the mixing develops inhomogeneity.
@@
    The turbulence is interaction-driven, i.e. transfer of excitation energy from neighboring cell-to-cell is comparable to, or exceeds local energy input from gradients. In this case, the production ratio for a given region $PR=\mathrm{\Delta}\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle/[-\int{\langle{\tilde{v}}_{r}\tilde{n}\rangle(\partial\langle n\rangle/\partial r)dr\ }]$ is $\geq$ 1. Here, $\mathrm{\Delta}\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle$ is the intensity flux differential across a region, and $\int{\langle{\tilde{v}}_{r}\tilde{n}\rangle(\partial\langle n\rangle/\partial r)dr\ }$ is the local production within it. $PR>1$ indicates that a region is driven primarily by interaction, i.e. by turbulence spreading or avalanching, and not by local instability. The production $PR$ is related to, but not identical to, the more familiar Kubo number $Ku\sim\tilde{v}\tau_{\mathrm{corr}}/\lambda_{\mathrm{corr}}$. $Ku\geq 1$ indicates correlated, non-diffusive mixing. Recent experiments have measured large PR and characterized its dependencies. Systems dominated by interactions tend to exhibit scale invariance across a range of scales bounded by the profile scale (large) and the cell size (small), so that transport phenomena exhibit a range of scales and so are necessarily inhomogeneous.
@@
    Coherent structure production. Strong local relaxation of gradients of a conserved order parameter (i.e. density, temperature) gives birth to a blob-void pair, i.e. see Figure 2. Blobs – particle excesses – propagate down gradient, while voids – particle deficits – propagate up gradient. These structures can enhance $PR$, excite waves and turbulence, and induce a range of spatial couplings. The effective structure propagation distance (“mean free path”) defines a new, emergent, length scale in the problem.
Figure 2: Schematic diagram of birth of a blob-void pair due to local relaxation of gradients of a conserved order parameter (density $n$ or temperature $T$).
All of process (i)-(iii) will tend to induce inhomogeneous mixing. Section 2 discusses the microphysics of mixing in plasma, including turbulence properties, coherent structures and their effects on entrainment, spreading and avalanching. Section 3 discusses the mesoscopic consequences of inhomogeneous mixing in plasma, namely $E\times B$ staircases. This section reports on a study of staircase structures and properties and presents results from a multi-machine comparison. Taken together, the results given in this paper paint a unique picture of the physics of layering in confined magnetized plasma as discussed in Section 4.
2 A close-up of inhomogeneous mixing: structures in turbulence
@@ 2.1 Basic characteristics of blobs/voids
In toroidal/linear magnetic confinement fusion edge plasmas of strongly turbulent regime, density or temperature fluctuations manifest intermittent/bursty behavior due to the presence of long-lived meso-scale coherent convective structures [1]. This kind of fluctuation structures/events, which exist within the ambient turbulence, are the so-called blobs and voids (or holes) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Intermittent density fluctuations have been extensively observed and investigated in magnetic confinement fusion, while there are only few reports on intermittent temperature fluctuations [13, 14]. Unless otherwise specified in the remainder of the article, blobs and voids refer to density blobs and density voids, respectively. The former are observed as magnetic-field-aligned filaments of excess density as compared with the background plasma, while the latter are observed as magnetic-field-aligned filaments of reduced density. Blobs and voids are born together to conserve density. In the two-dimensional cross-section normal to the magnetic field, these structures look just like “blobs” and “voids” [3, 6, 15, 16]. Blobs and void structures can be intuitively reflected through the two-dimensional cross-conditional averages [17] of ion saturation current $\tilde{I}_{sat}$ (representative of density fluctuations), as shown by Figures 3(c) and 3(d), respectively [15].
@@
Figure 3: Blob and void structures in the linear device LAPD [15]. In (a) and (b) are cross-conditional averages of ion saturation current $\tilde{I}_{sat}$ (representative of density fluctuations) for blob and void events respectively, showing propagation of blobs out of the plasma and voids back into the plasma. In (c) and (d) are two-dimensional cross-conditional averages to show the blob and void structures respectively. Here, a negative $\Delta x$ indicates closer to the outer wall, and a positive $\Delta x$ indicates closer to the core of plasma. Reprinted from Carter et al. [15]. Copyright 2006 AIP.
@@
Experimental studies show that blobs and voids can be observed at the boundary region of magnetic confinement plasma, either inside the last closed flux surface (LCFS) or outside the LCFS (called scrape-off layer, SOL) on the low field side [11, 18, 19, 20]. The positive/negative skewness of density fluctuations has been widely used to identify the local predominance of blobs/holes, respectively [1, 5]. Theoretical work indicates that the origin of blobs and voids is due to instabilities that exhibit strong intensity on the low field side with bad curvature, such as curvature-driven interchange instability, drift wave, etc [21, 22, 23]. The $E\times B$ sheared flow is also an important parameter controlling the blobs/voids’ generation [24, 25]. These explain why the blobs and voids are detected at the edge location of maximum normalized pressure gradient or inside the edge shear layer [6, 26]. In the presence of a charge-dependent drift, e.g., curvature drift and grad-B drift, charge separation and the resulting $E\times B$ drift cause the blobs to move outward to the outer wall (down the mean gradient), while the voids propagate inward to the core (up the mean gradient) [2]. The opposite propagation directions of blobs and voids can be apparently seen from the opposite time delays of cross-conditional averages of density fluctuations for blob and void events at different radial locations as shown in Figures 3(a) and 3(b), respectively.
@@
Blobs and voids contribute significantly to the radial transport of energy and particle across the edge and scrape-off layer (SOL) [27, 28, 29, 30, 31]. It should be noted that the core-boundary coupling is a key issue for future nuclear fusion reactors such as ITER and DEMO, that is, simultaneously achieving high-performance core plasma (to obtain high fusion power) and high-dissipation boundary plasma (to protect the plasma facing materials) [32, 33, 34]. Therefore, it is necessary to study blobs and voids that are active in the plasma edge layers between the core and boundary. In recent years, new experimental research advancements and discoveries have provided novel insights and perspectives to understand the physical behavior and critical influence of blobs/voids. These include: (i) turbulence spreading induced by blobs/voids; (ii) turbulent layer broadened by blobs/voids; and (iii) interactions between blobs/voids and sheared flow, which will be elaborated in detail in the following parts.
@@ 2.2 Turbulence spreading induced by blobs/voids
Turbulence spreading means the propagation of turbulence intensity or energy in space. This is often quantified as the flux of turbulence intensity or energy across the magnetic field. For example, the turbulence spreading flux for density turbulence intensity is calculated as $\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle/2$ [23, 35, 36]. Spreading of turbulence occurs via relaxation of inhomogeneous turbulence and entrainment of laminar or more weakly turbulent regions [37]. One well known example is an expanding wake formed downstream of a moving object, which corresponds to turbulence spreading from the fully turbulent core [38]. The turbulence intensity field can be decoupled from the local instability growth rate due to turbulence spreading. An in-depth experimental research of the relation of turbulence spreading and blobs-voids in edge plasmas via Langmuir probe measurements on J-TEXT tokamak has been reported recently [7]. It is found that the turbulence spreading flux originated from positive density fluctuations is positive, meaning the outward propagation of the positive $\tilde{n}$ turbulence intensity. The turbulence spreading flux originated from negative density fluctuations is negative, meaning the inward propagation of the negative $\tilde{n}$ turbulence intensity. This is shown by the turbulence spreading “lip” curves in Figure 4(a), which presents the distribution of turbulence spreading flux $P_{spreading}(A)=\sum_{A-0.05<|\tilde{n}|/\sigma_{\tilde{n}}<A+0.05}(\tilde{v}_{r}\tilde{n}^{2}/2)/M$ for each positive and negative $\tilde{n}$, respectively [7]. $M=6000$ indicates the total number of sampling points of $\tilde{n}$ and $\tilde{v}_{r}$. The distribution curve of turbulent spreading flux originated from positive density fluctuations is always positive (corresponding to outward spreading), i.e. $P_{spreading}(A>0)>0$, which outlines the upper edge of the “lip”. The distribution curve of turbulent spreading flux originated from negative density fluctuations is always negative (corresponding to inward spreading), i.e. $P_{spreading}(A<0)<0$, which outlines the lower edge of the lip. The inward turbulence spreading caused by voids has also been further demonstrated by the experimental measurements via beam emission spectroscopy (BES) on DIII-D tokamak [39] and is the subject of a newly developed theoretical model [9].
@@
[…]
@@
The result above highlights the importance of blobs/voids convective structures in conveying the spreading of turbulence in the edge plasmas, and brings about a challenge to diffusive models of turbulence spreading in the form of Fick’s law as often assumed and long used [7, 36]. Correlatively, experimental observations using electron cyclotron emission (ECE) diagnostic on tokamak plasmas indicate that non-diffusive avalanche-like transport events, composed of temperature voids and temperature blobs propagating inwards and outwards, respectively, prevail in the electron heat transport channel in core plasmas [40, 41, 42, 43].
2.3 Turbulent layer broadened by blobs/voids

### v2

@@ Abstract
Inhomogeneous mixing and the consequent mesoscopic layered structure have been observed in many physical systems, including magnetically confined fusion plasmas. Especially, in plasmas, mixing can be enhanced through turbulence spreading by intermittent coherent structures (blobs/voids), or suppressed due to the formation of transport barriers (sheared zonal flows). Interestingly, blobs/voids and zonal flows are not independent, and they can co-exist in a state of inhomogeneous mixing, often called the $\mathbf{E}\times\mathbf{B}$ staircase. In this paper, we first introduce recent experimental progress on the physics of blobs/voids: how turbulence spreading by blobs/voids occurs, the consequences of enhanced turbulence spreading for the scrape-off layer (SOL) power decay length, and the interaction between blobs/voids and zonal flows. Then, we provide a brief review of experimental results on staircases, or more generally layered mesoscopic transport barriers. Staircases are often elusive in experiments, requiring integrating multi-diagnostic data, utilizing high-dimensional diagnostics, or extracting hidden information from signals for their identification. Our understanding is still incomplete. This paper serves as an initial step toward applying insights gained from inhomogeneous mixing due to blobs/voids to the understanding of a staircase.
@@ 1 Introduction
Inhomogeneous mixing is a transport or relaxation process which is not uniform. So, in Figure 1, case (a) depicts homogeneous mixing, while case (b) shows inhomogeneous mixing. Note there, the mixing zones don’t overlap. As suggested by Figure 1, inhomogeneous mixing of a profile will quite naturally produce layered or staircase structures on mesoscales or macroscales. Such layered structure is the focus of this collection, and staircases in magnetized plasma are the foci of this paper. Inhomogeneous mixing and staircases form in many physical systems, including saltwater via thermohaline convection, stirred stratified fluids, geophysical fluids (i.e. potential vorticity (PV) staircases), Cahn-Hilliard Navier-Stokes fluids via phase separation and magnetized plasma, the subject of this volume. Inhomogeneous mixing has a bi-stable or multi-stable, and multi-scale character, which is intrinsic to the macroscopic layering, i.e. different regions mix unequally. This paper discusses inhomogeneous mixing and staircase structure in confined magnetized plasma.
Figure 1: Toroidal plasma confined by toroidal and poloidal magnetic fields. Plasma is heated and fueled by external devices, forming the peaked density/temperature profile for fusion reactions. Plasma instabilities, growing due to the gradient of profiles, can develop into plasma turbulence which mixes density/temperature. Schematic diagram of (a) homogeneous mixing and (b) inhomogeneous mixing of density/temperature. Panel (b) shows a staircase in plasma density/temperature, with flat regions representing well-mixed regions and jumps representing transport barriers to mixing.
In tokamaks or stellarators, plasma is confined within nested magnetic flux surfaces formed by magnetic fields. The field lines lie on the flux surfaces. The geometric structure of the field lines is characterized by the safety factor $q\approx\frac{rB_{\phi}}{RB_{\theta}}$ where $r$ and $R$ are minor and major radius and $B_{\phi}$ and $B_{\theta}$ are toroidal and poloidal magnetic field strength, respectively (Figure 1). It represents the ratio between the number of toroidal rotations and the number of poloidal rotations as one follows a field line. The flux surface of the rational $q$ value is called the rational surface on which the field lines are closed lines. For example, on the flux surface of $q=3/1$, a field line starting from any point will eventually return to its original position after completing three toroidal rotations and one poloidal rotation.
Turbulence in plasma (at low $\beta=(\mathrm{thermal~plasma~pressure/magnetic~pressure})$) consists of an ensemble of spatially localized convective cells. They are often shaped to have the strong anisotropy ($k_{\parallel}/k_{\perp}\ll 1$ where $k_{\parallel}$ and $k_{\perp}$ are the parallel and perpendicular components of the wavenumber vector $\mathbf{k}$ with respect to the magnetic field $\mathbf{B}$) and are pinned to sites determined by the confining magnetic geometry. $\mathbf{k}\cdot\mathbf{B}=0$ resonances or rational surfaces are one such pinning site. Typically, cells have radial correlation length $\lambda_{\mathrm{corr}}$ that is $\lambda_{\mathrm{corr}}/L_{p}\ll 1$. Here, $L_{p}=\frac{\langle p\rangle}{d\langle p\rangle/dr}$ is a plasma (density or temperature) profile scale length. At the edge, $\lambda_{\mathrm{corr}}/L_{p}\leq 1$. There are three obvious means by which mixing by plasma turbulence develops inhomogeneity.
@@
    Along with particles, heat, etc., the turbulence mixes plasma potential vorticity – and thus vorticity – so $\langle{\tilde{v}}_{r}\nabla^{2}\tilde{\phi}\rangle\neq 0$ where $\tilde{v}_{r}$ and $\tilde{\phi}$ are fluctuations of radial plasma velocity and plasma potential, respectively. Here, $\langle\rangle$ means averaging on the flux surface. The Taylor identity then gives $\langle{\tilde{v}}_{r}\nabla^{2}\tilde{\phi}\rangle=-\partial\langle{\tilde{v}}_{r}{\tilde{v}}_{\theta}\rangle/\partial r$ where $\tilde{v}_{\theta}$ is poloidal velocity fluctuation, so a non-zero Reynolds force develops, and drives sheared zonal flows on the scale of the spectral envelope. These zonal flows sink energy from the underlying relaxation process and thus impose a variable shearing pattern on the turbulence. Consequently, mixing develops inhomogeneity.
@@
    The turbulence is interaction-driven, i.e. transfer of excitation energy from neighboring cells is comparable to, or exceeds, local energy input from gradients. In this case, the production ratio for a given region, which can be written as $PR=d(\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle)/[-\int{\langle{\tilde{v}}_{r}\tilde{n}\rangle(\partial\langle n\rangle/\partial r)dr\ }]$, is $\geq$ 1. Here, $\tilde{n}$ is plasma density fluctuation, $d(\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle)$ is the intensity flux differential across a region, and $\int{\langle{\tilde{v}}_{r}\tilde{n}\rangle(\partial\langle n\rangle/\partial r)dr\ }$ is the local production within it. $PR>1$ indicates that a region is driven primarily by interaction, i.e. by turbulence spreading or avalanching from other regions, and not by local instability. Note that above definition of the production ratio is relevant to edge turbulence. Regarding core plasmas, ion temperature gradient (ITG) turbulence is a natural system for which to calculate a production ratio (a brief derivation of a production ratio for the gyrokinetic ITG model is given in Appendix for interested readers). $PR$ is related to, but not identical to, the more familiar Kubo number $Ku\sim\tilde{v}\tau_{\mathrm{corr}}/\lambda_{\mathrm{corr}}$ where $\tau_{\mathrm{corr}}$ is the cell correlation time. $Ku\geq 1$ indicates correlated, non-diffusive mixing. Recent experiments have measured large $PR$ and characterized its dependencies. Systems dominated by interactions tend to exhibit scale invariance across a range of mesoscopic scales $l$ bounded by the mean profile scale length ($L_{p}$) and the cell correlation length ($\lambda_{\mathrm{corr}}$), i.e. $\lambda_{\mathrm{corr}}<l<L_{p}$. Thus, transport phenomena can manifest on a broad range of scales, so the mixing likely is inhomogeneous.
@@
    Coherent structure production. Strong local relaxation of gradients of a conserved order parameter (i.e. density or temperature) gives birth to a blob-void pair (see Figure 2). Blobs – particle excesses – propagate down-gradient, while voids – particle deficits – propagate up-gradient. These structures can enhance $PR$, excite waves and turbulence, and induce a range of spatial couplings. The effective structure propagation distance (“mean free path”) defines a new, emergent, length scale in the problem.
Figure 2: Schematic diagram of birth of a blob-void pair due to local relaxation of gradients of a conserved plasma density $n$ or temperature $T$.
[…]
2 Close-up of inhomogeneous mixing: blobs/voids in turbulence
@@ 2.1 Basic characteristics of blobs/voids
In the strongly turbulent edge region of toroidal/linear magnetically confined fusion plasmas, density or temperature fluctuations manifest intermittent/bursty behavior due to the presence of long-lived mesoscale coherent convective structures [1]. This kind of fluctuation structures/events, which exist within the ambient turbulence, are the so-called blobs and voids (or holes) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Intermittent density fluctuations have been extensively observed and investigated in magnetic confinement fusion, while there are only few reports on intermittent temperature fluctuations [13, 14]. Unless otherwise specified in the remainder of the article, blobs and voids refer to density blobs and density voids, respectively. The former are observed as magnetic-field-aligned filaments of excess density as compared with the background plasma, while the latter are observed as magnetic-field-aligned filaments of reduced density. Blobs and voids are born together to conserve density. In the two-dimensional cross-section normal to the magnetic field, these structures look just like “blobs” and “voids” [3, 6, 15, 16]. Blob and void structures can be intuitively identified through the two-dimensional cross-conditional averages [17] of ion saturation current $\tilde{I}_{sat}$ (representative of density fluctuations), as shown by Figures 3(c) and 3(d), respectively [15].
@@
Figure 3: Blob and void structures in the linear device LAPD [15]. In (a) and (b) are cross-conditional averages (CA) of ion saturation current $\tilde{I}_{sat}$ (representative of density fluctuations) for blob and void (hole) events, respectively, showing propagation of blob out of the plasma and void back into the plasma. Black line represents CA of $\tilde{I}_{sat}$ measurements by the reference (event detection) channel, and other lines represent CA of $\tilde{I}_{sat}$ measurements by channels at radially separated positions. For example, the red line is CA of $\tilde{I}_{sat}$ measurements by the channel which is 2 cm inward (closer to the plasma center) from the reference channel which is used to detect blob/void events. In (c) and (d) are two-dimensional cross-conditional averages to show the blob and void structures, respectively. $\Delta x$ and $\Delta y$ mean the relative measurement position against the reference channel position in the radial and poloidal direction, respectively. Reprinted from Carter et al. [15]. Copyright 2006 AIP.
@@
Experimental studies show that blobs and voids can be observed at the boundary region of magnetic confinement plasma, either inside the last closed flux surface (LCFS) or outside the LCFS (in the scrape-off layer, SOL) on the low field side [11, 18, 19, 20]. The positive/negative skewness of density fluctuations has been widely used to identify the local predominance of blobs/holes, respectively [1, 5]. Theoretical work indicates that the origin of blobs and voids is caused by instabilities such as curvature-driven interchange instability and drift wave instability [21, 22, 23]. The former exhibits a strong intensity on the low magnetic field side with bad curvature in toroidal plasmas, while the latter can exist both in toroidal and linear plasmas. The $\mathbf{E}\times\mathbf{B}$ sheared flow is also an important parameter controlling the blobs/voids’ generation [24, 25]. These explain why the blobs and voids are detected at the edge location of maximum normalized pressure gradient or inside the edge shear layer [6, 26]. In the presence of a charge-dependent drift, e.g., curvature drift and grad-B drift, charge separation and the resulting $\mathbf{E}\times\mathbf{B}$ drift cause the blobs to move outward to the outer wall (down the mean gradient), while the voids propagate inward to the core (up the mean gradient) [2]. The opposite propagation directions of blobs and voids can be seen from the opposite time delays of cross-conditional averages (CA) of density fluctuations for blob and void events at different radial locations as shown in Figures 3(a) and 3(b), respectively.
@@
Blobs and voids contribute significantly to the radial transport of energy and particle across the edge and SOL regions [27, 28, 29, 30, 31]. It should be noted that the core-boundary coupling is a key issue for future nuclear fusion reactors such as ITER and DEMO, that is, simultaneously achieving high-performance core plasma (to obtain high fusion power) and high-dissipation boundary plasma (to protect the plasma facing materials) [32, 33, 34]. Therefore, it is necessary to study blobs and voids that are active in the plasma edge layers between the core and boundary. In recent years, new experimental research advancements and discoveries have provided novel insights and perspectives to understand the physical behavior and critical influence of blobs/voids. These include: (i) turbulence spreading induced by blobs/voids; (ii) turbulent layer broadening by blobs/voids; and (iii) interactions between blobs/voids and sheared flow, which will be discussed in detail below.
@@ 2.2 Turbulence spreading induced by blobs/voids
Turbulence spreading means the propagation of turbulence intensity or energy in space. This is often quantified as the flux of turbulence intensity or energy across the magnetic field. For example, the turbulence spreading flux for density turbulence intensity is calculated as $\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle/2$ [23, 35, 36]. Spreading of turbulence occurs via relaxation of inhomogeneous turbulence and entrainment of laminar, or more weakly turbulent regions [37]. One well known example is an expanding wake formed downstream of a moving object, which corresponds to turbulence spreading from the fully turbulent core [38]. The turbulence intensity field can be decoupled from the local instability growth rate due to turbulence spreading. In-depth experimental research of the relation of turbulence spreading to blobs-voids in edge plasmas via Langmuir probe measurements on J-TEXT tokamak has been reported recently [7]. It is found that the turbulence spreading flux originating from positive density fluctuations is positive, meaning the outward propagation of the positive $\tilde{n}$ turbulence intensity. The turbulence spreading flux originating from negative density fluctuations is negative, meaning the inward propagation of the negative $\tilde{n}$ turbulence intensity. This is shown by the turbulence spreading “lip” curves in Figure 4(a), which presents the distribution of turbulence spreading flux $P_{spreading}(A)=\sum_{A-0.05<|\tilde{n}|/\sigma_{\tilde{n}}<A+0.05}(\tilde{v}_{r}\tilde{n}^{2}/2)/M$ calculated using either positive or negative $\tilde{n}$ where $A=|\tilde{n}|/\sigma_{\tilde{n}}$ and $\sigma_{\tilde{n}}$ is the standard deviation of plasma density fluctuation [7]. $M=6000$ indicates the total number of sampling points of $\tilde{n}$ and $\tilde{v}_{r}$. The distribution curve of turbulent spreading flux originating from positive density fluctuations is always positive (corresponding to outward spreading), i.e. $P_{spreading}(A)>0$ for $\tilde{n}/\sigma_{\tilde{n}}>0$, which outlines the upper edge of the “lip”. The distribution curve of turbulent spreading flux originating from negative density fluctuations is always negative (corresponding to inward spreading), i.e. $P_{spreading}(A)<0$ for $\tilde{n}/\sigma_{\tilde{n}}<0$, which outlines the lower edge of the lip. The inward turbulence spreading caused by voids has also been further demonstrated by the experimental measurements via beam emission spectroscopy (BES) on the DIII-D tokamak [39] and is the subject of a newly developed theoretical model [9].
@@
The net total turbulence spreading flux $\langle{\tilde{v}}_{r}{\tilde{n}}^{2}\rangle/2$ is taken as the sum of positive and negative components, and the weak, moderate, and strong spreading flux cases are shown in Figure 4(a). The distribution function of turbulence spreading flux is almost symmetric for the weak spreading case (red diamonds). The distribution function becomes more strongly asymmetric (green crosses and blue inverted triangles) for the stronger net outward spreading flux cases. As shown in Figure 4(b), while the contribution by the negative high-amplitude density fluctuations to the net flux decreases, the contribution by the positive high-amplitude density fluctuations increases, which is especially true for contributions with absolute amplitudes higher than $2\sigma_{\tilde{n}}$. The outward turbulence spreading is caused by the blobs-dominant events with $\tilde{n}\geq 2\sigma_{\tilde{n}}$, and the inward spreading is caused by the holes-dominant events with $\tilde{n}\leq-2\sigma_{\tilde{n}}$. This blob-void asymmetry emerges as crucial to turbulence spreading.
@@
The result above highlights the importance of blob/void convective structures in conveying the spreading of turbulence in the edge plasmas, and brings about a challenge to diffusive models of turbulence spreading in the form of Fick’s law as often assumed and long used [7, 36]. Also, turbulence spreading is closely related to avalanche transport [36] in core plasmas. The mechanism of turbulence spreading described above is expected to be critical in understanding the avalanche transport phenomena observed in core plasmas [40, 41, 42, 43, 44] as well.
2.3 Turbulent layer broadening by blobs/voids